OpenStudy (anonymous):

Marta is making a scale drawing of her bedroom. Her rectangular room is 12.5 feet wide and 15 feet long. On the scale drawing, the width of her room is 5 inches. What is the length on her drawing?

OpenStudy (lgbasallote):

use ratio and proportion here. since you're looking for length let that equal x 12.5 : 15 = 5 : x \[\frac{12.5}{15} = \frac{5}{x}\] \[12.5 x = 5 \times 15\] \[x = \frac{5 \times 15}{12.5}\]